# Understanding uncertainty in Bayesian cluster analysis

### v1

@@ 4 Experiments
We now consider several synthetically generated datasets to illustrate scenarios where WASABI can be beneficial in explaining the uncertainty in the posterior distribution. We consider two main scenarios where the posterior distribution displays high uncertainty: (1) when the model is correctly specified, but the mixture components are not well-separated, and (2) when the model is misspecified.
@@
First, we consider scenario (1), by examining different levels of cluster separation in Example 2. Specifically, several values of $m$, defining the means (±$(\pm m,\pm m)$ of the four mixture components, are considered (see Figure 11 in the Appendix). For each value of $m$, we generate 10 datasets. For small values of $m$, the four mixture components overlap, making the distribution concentrated around the partition with one cluster. In contrast, when $m$ is large, the four components are well separated and easy to distinguish.
@@
To quantify the uncertainty in the posterior, we can measure the posterior distribution’s spread around its center by the EVI. This is displayed in the left panel of Figure 4, showing how the posterior uncertainty starts low and increases when there is a moderate level of cluster separation, and then decreases when the cluster separation is high. The right panel of Figure 4 instead shows the improvement in the Wasserstein distance achieved by WASABI, quantified by $\textrm{EVI}(\rho^{*})-W_{\text{VI}}(\hat{\pi},q^{*})$, when summarizing with $L=10$ particles instead of a single estimator. These figures show that WASABI can explain the variation in the posterior the most when the posterior uncertainty is highest (i.e. for values of $m\approx 1.3$).
@@
Figure 4: Experiment with different levels of cluster separation, determined by $m$. Left panel: posterior spread around the minVI estimator, quantified by its expected VI. Right panel: improvement in Wasserstein distance achieved by WASABI for different values of $m$; the points are colored by the number of clusters in the corresponding minVI estimate.
@@
In scenario (2), we explore several examples of misspecified models with data generated from different distributions: (a) a bivariate Gaussian distribution truncated to the unit square, and (b) a bivariate skewed-t distribution (assuming independence). To investigate the effect of misspecification, in (a), the standard deviation is varied, so that small values are well-specified and larger values resemble a uniform distribution on the unit square; and in (b), the degrees of freedom and skewness parameter are varied. A detailed description of these synthetic experiments can be found in Section D of the Appendix.
@@
Figure 5: Synthetic experiment with misspecified models. The top two rows show results for truncated Gaussian and skewed-t ($df=5$) distributions. Left: improvement in Wasserstein distance by WASABI versus posterior spread (EVI); central/right: particle scatterplots with cluster assignments and a summary plot, for one synthetic dataset. The bottom row shows Wasserstein improvement for skewed-t with varying degrees of freedom.
@@
Figure 5 displays the improvement in Wasserstein distance achieved by WASABI as a function of the posterior spread (measured by EVI) for (a) the truncated Gaussian distribution (top row, left panel), and (b) for the skewed-t distribution with $df=5$ (second row, left panel). As expected, overall WASABI tends to explain higher variations in the posterior when the EVI is larger. This also correlates with increasing misspecification; for the truncated Gaussian distribution (a), the EVI increases with the standard deviation, and for the skewed-t distribution with $df=5$ (b), larger values of the skewness parameter result in higher EVI and larger improvement by WASABI. The same improvement is visible in the bottom row (all panels) for the skewed-t distribution with varying degrees of freedom, where we can also observe that most of the variation is driven by the skewness parameter, but for smaller degrees of freedom, higher posterior uncertainty (and improvement achieved by WASABI) and observed. Finally, the central and right panels in the first two rows provide a summary of the data and particles for one representative example, displaying median values of EVI and improvement in Wasserstein distance; for the truncated Gaussian, this corresponds to $\sigma=0.4$, while for the skewed-t distribution it corresponds to $df=5$ and skewness parameter equal 2.
@@ 5 Real data analysis
The human papillomavirus (HPV) vaccine, a key tool in reducing cervical and other cancers, is routinely offered to 14-year-old girls in Scotland. However, decline in uptake is a global concern, underscoring the need to examine factors that influence this pattern. Here, we focus on studying how neighborhood deprivation may affect HPV vaccine uptake in Scotland. The data [^2] Data available through Public Health Scotland: https://scotland.shinyapps.io/ScotPHO_profiles_tool/ consists of the percentage of HPV uptake, denoted by $y_{i}$, and the deprivation level, denoted by $x_{i}$ and measured as the proportion of young people living in the most income deprived quintile, across the $n=1279$ Scottish Intermediate Zones. For a visualization of HPV uptake across these zones, see the left panel of Figure 6.
@@ 6 Discussion
Understanding the uncertainty characterized by the posterior distribution is challenging in complex, high-dimensional Bayesian models. Additionally, as the posterior is intractable in such settings, algorithms for approximate inference, such as MCMC, are employed, which can return tens to hundreds of thousands of approximate posterior draws. In this work, we focus on clustering, where the posterior is defined on the high-dimensional, partially-ordered space of partitions. In this non-standard setting, we propose a WASserstein Approximation for Bayesian cluserIng (WASABI) to summarize the posterior draws with a small number of weighted partition estimates. Through extensive simulations and analyses, we showcase its utility in practical settings when some clusters are not well separated or the model may be misspecified. Indeed, while many applications focus on a single clustering estimate, we demonstrate how it can be misleading in such settings and how WASABI captures the different modes of clustering.
The Wasserstein approximation requires selecting a metric on the partition space. We focus on the variation of information (Meilă, 2007), also exploiting its properties to propose novel tools to assess uncertainty in the cluster assignment of individual points. However, this can be easily extended to other choices, such as Binder’s loss, ARI, or generalized versions (Dahl et al., 2022), or to more general settings, e.g. Buch et al. (2024) develop a Bayesian approach to clustering in the presence of noise.
An efficient algorithm, along with various tools to visualize and quantify uncertainty, is provided in the R package at https://github.com/cecilia-balocchi/WASABI. In practice, only a small number of iterations are typically required, with each one scaling linearly with the number of data points and MCMC samples. To reduce the cost of the latter, WASABI allows subsampling of the MCMC draws. For the former, future work will explore combining WASABI with distributed methods for Bayesian clustering (Ni et al., 2020).
WASABI is useful for improved understanding of the posterior, but could also be relevant for posterior compression in order to reduce both storage (saving only the WASABI particles and weights) and the computational cost of predictive inference. For example, density estimation can be expensive, depending on the size of the grid or number of new locations. For Example 1, density estimates are provided in Section D.1 of the Appendix, where the WASABI-based estimates significantly reduce the cost by a factor of $1,000$ compared to MCMC, at only a small increase in error.
Beyond clustering, the Wasserstein approximation could be beneficial for other Bayesian models. Future work, will investigate settings, such as variable selection, trees, networks, and more.
@@ Appendix A Proof to Proposition 1
Let $\bm{w}^{*}=(w^{*}_{1},\ldots,w^{*}_{L})$ and $\bm{\rho}^{*}=\{\rho^{*}_{1},\ldots,\rho^{*}_{L}\}$ respectively denote the vector of weights and the set of atoms of $q^{*}(\cdot)=\sum_{\ell=1}^{L}w^{*}_{\ell}\delta_{\rho^{*}_{\ell}}(\cdot)$. To solve (2) we need to find the optimal $\bm{w}^{*}$ and $\bm{\rho}^{*}$, but also for a given $q$ we need to find the optimal coupling $J^{*}(q)$ that minimizes $\sum_{\rho,\rho^{\prime}}d_{VI}(\rho,\rho^{\prime})J(\rho,\rho^{\prime})$ and that satisfies the marginal constraints.
@@ Appendix B Detailed algorithm and initialization strategies
  • Average Linkage Initialization: A hierarchical clustering is performed on the posterior similarity matrix using average linkage, and the $L$ resulting partitions with number of clusters ranging from $1$ to $L$ (from appropriate tree cuts) are selected as initial centers. This approach specifies the initialization with diverse partitions that represent stable co-clustering patterns in the MCMC sample.
  • Complete Linkage Initialization: Similarly, hierarchical clustering using complete linkage is applied to the posterior similarity matrix. The dendrogram is cut at levels corresponding to $1$ up to $L$ clusters, and the resulting $L$ partitions are used as the initial medoids. Complete linkage generally favors compact, small-diameter clusters.
  • Fixed Initialization: This strategy allows the user to directly specify a set of $L$ fixed partitions to be used as initial centers. This is useful, for example, if the user wishes to include certain known clusterings (such as the minVI estimator, MAP clustering, or expert-proposed solutions) as initial candidates. Alternatively, the EVI can be computed for all MCMC samples, and the $L$ partitions with the smallest values are passed as starting points.
  • k-means++ style Initialization (+++++ +): Analogous to the k-means++ technique (Vassilvitskii and Arthur, 2006), this strategy promotes diversity among initial centers by iteratively choosing the next center with probability proportional to its VI distance from the closest already chosen center. This increases the likelihood that the $L$ initial medoids will be well separated in partition space, which often leads to better local optima.
  • Random Partition Initialization: In this approach, each data point is randomly assigned to one of $L$ groups, and the center for each group is chosen based on these assignments. This tends to produce initial medoids that are close together in partition space, often near the “center” of the sample distribution. Although not suited for finding very distinct particles, it may be useful for robustness checks.
  • Enhanced Initialization (+++++++ + +): This extends the +++++ + initialization by running the +++++ + initialization search on the set of the MCMC samples combined with the partitions obtained from the average and complete linkage strategies. The intention is to expand the diversity of the initial set and include particularly informative candidates, possibly leading to a broader exploration of the solution space.
  • Hierarchical Super-Initialization (topvi): In this method, the expected VI is computed for all the partitions generated by the average linkage, complete linkage, and (if supplied) fixed initialization methods. The $L$ best partitions among these (with lower EVI) are then selected as initial centers. This approach mimics the idea of selecting the partitions in the MCMC sample with lowest EVI, while avoiding the computational complexity of computing the EVI for all MCMC samples.
Note: When practical, it is helpful to combine several strategies and perform multiple runs with each, then select the solution yielding the lowest objective (Wasserstein) distance. In particular, using k-means++ or hierarchical strategies offers a good trade-off between computational cost and solution diversity, while random initialization is useful for unbiased robustness assessment.
@@
         Update centers: $\rho^{*}_{\ell}\leftarrow\textsc{minVI}(\mathcal{N}_{\ell})$triangleright Different search-algorithms can be used
@@
[…]
@@ Appendix C Additional material on describing and visualizing the particles
In Section 3.2, we introduce Example 2, where the data is generated from four Gaussian mixtures. We fit a diagonal location-scale Dirichlet Process Mixture model, using the R package BNPmix (Corradin et al., 2021), run for 1,000 iterations after a burn-in of 10,000 iterations. In the posterior similarity matrix shown in Figure 7, uncertainty about the number of clusters is apparent, but a subtle pattern of four equal-sized blocks can be observed, even though some of these seem to be often merged together.
@@
Figure 8 displays the posterior similarity matrices (PSMs) for the regions of attraction associated with each WASABI particle in Example 2. While Figure 7 (in Section 3.2) presents PSMs labeled by each particle’s own clusters (providing insight into the uncertainty structure around each specific clustering) here, the rows and columns are labeled according to the clusters defined by the meet of all particles. This common labeling facilitates direct comparison of cluster uncertainty across different regions of attraction, helping to identify points or groups whose allocation is most variable or most consistent between the particles represented by the WASABI summary.

### v2

@@ 4 Experiments
We now consider several simulated datasets to illustrate scenarios when WASABI can be beneficial in explaining the uncertainty in the posterior. We consider two main settings: (1) the model is correctly specified, but the mixture components are not well-separated, and (2) the model is misspecified.
@@
First, we consider scenario (1), by examining different levels of cluster separation in Example 2. Specifically, several values of $m$, defining the means $(\pm m,\pm m)$ of the four mixture components, are considered (see Figure 14 in the Appendix). For each value of $m$, we generate 10 datasets. When $m$ is small, the four components overlap, making the posterior concentrated around the partition with one cluster. Instead, when $m$ is large, the four components are well separated and easy to distinguish.
@@
To quantify the uncertainty in the posterior, we measure the spread around its center by the EVI. This is displayed in the left panel of Figure 4, showing how the posterior uncertainty starts low and increases when there is a moderate level of cluster separation, and then decreases when the cluster separation is high. The right panel of Figure 4 instead shows the improvement in the Wasserstein distance achieved by WASABI, quantified by $\textrm{EVI}(\rho^{*})-W_{\text{VI}}(\hat{\pi},q^{*})$, when summarizing with $L=10$ particles instead of a single estimator. These figures show that WASABI explains the variation in the posterior the most when the posterior uncertainty is highest (i.e. for values of $m\approx 1.3$).
@@
Figure 4: Experiment with different levels of cluster separation, determined by $m$. Left panel: posterior spread around the minVI estimator, quantified by its EVI. Right panel: improvement in Wasserstein distance achieved by WASABI for different values of $m$; the points are colored by the number of clusters in the corresponding minVI estimate.
@@
In scenario (2), we explore several examples of misspecified models with data generated from different distributions: (a) a bivariate Gaussian distribution truncated to the unit square, and (b) a bivariate skewed-t distribution (assuming independence). To investigate the effect of misspecification, in (a), the standard deviation is varied, so that small values are well-specified and larger values resemble a uniform distribution on the unit square; and in (b), the degrees of freedom and skewness parameter are varied. A detailed description of these experiments is in Section D.2 of the Appendix.
@@
Figure 5: Experiment with misspecified models. The top two rows show results for truncated Gaussian and skewed-t ($df=5$) distributions. Left: improvement in Wasserstein distance by WASABI versus posterior spread (EVI); central/right: particle scatterplots with cluster assignments and a summary plot, for one simulated dataset. The bottom row shows Wasserstein improvement for the skewed-t with varying degrees of freedom.
@@
Figure 5 displays the improvement in Wasserstein distance achieved by WASABI as a function of the posterior spread (measured by EVI) for (a) the truncated Gaussian distribution (top row, left panel), and (b) for the skewed-t distribution with $df=5$ (second row, left panel). Overall, WASABI tends to explain higher variations in the posterior when the EVI is larger. This also correlates with increasing misspecification; for the truncated Gaussian distribution (a), the EVI increases with the standard deviation, and for the skewed-t distribution with $df=5$ (b), larger values of the skewness parameter result in higher EVI. The same improvement is visible in the bottom row (all panels) for the skewed-t distribution with varying degrees of freedom, where we also observe that most of the variation is driven by the skewness parameter, but for smaller degrees of freedom, higher posterior uncertainty (and improvement by WASABI) is observed. Finally, the central and right panels in the first two rows illustrate the WASABI results for one example, with median EVI value, corresponding to $\sigma=0.4$ for the truncated Gaussian, and to $df=5$ and skewness of 2 for the skewed-t.
@@ 5 Real data analysis
5.1 Effects of neighborhood deprivation on HPV vaccine uptake
The human papillomavirus (HPV) vaccine, a key tool in reducing cervical and other cancers, is routinely offered to 14-year-old girls in Scotland. However, decline in uptake is a global concern, underscoring the need to examine factors that influence this pattern. Here, we focus on studying how neighborhood deprivation may affect HPV vaccine uptake in Scotland. The data [^1] Data available through Public Health Scotland: https://scotland.shinyapps.io/ScotPHO_profiles_tool/ consists of the percentage of HPV uptake, denoted by $y_{i}$, and the deprivation level, denoted by $x_{i}$ and measured as the proportion of young people living in the most income deprived quintile, across the $n=1279$ Scottish Intermediate Zones. The left panel of Figure 6 visualizes HPV uptake across these zones.
@@ 5.1 Effects of neighborhood deprivation on HPV vaccine uptake
5.2 Investigating projection patterns at the single-neuron resolution
Mapping long-range connectivity in the brain is crucial to establish biological mechanisms of cognition and understand brain dysfunction in disorders. Recent technology (e.g. Kebschull et al., 2016; Chen et al., 2019) allows mapping axonal projections at the level of individual neurons by tagging thousands of neurons in parallel with unique mRNA barcodes and employing sequencing of tissues biopsies at target areas. Since the RNA is transported to the axon terminals, the observed barcode counts for each neuron reflect the relative strength of projection to the target areas. Compared with bulk-labeling technologies, this has revealed unexpected diversity in projection patterns among otherwise similar neurons. To disentangle this heterogeneity, a novel statistical method for barcode counts (HBMAP, Wade et al., 2025) employs hierarchical Bayesian mixtures to identify projection patterns and integrate data across multiple experiments.
We consider data from Chen et al. (2019) measuring projections from the auditory cortex to the 11 target regions (the orbitofrontal cortex (OFC), motor cortex (Motor), rostral striatum (Rstr), somatosensory cortex (SSctx), caudal striatum (Cstr), amygdala (Amyg), ipsilateral visual cortex (VisIp), contralateral visual cortex (VisC), contralateral auditory cortex (AudC), thalamus (Thal), and tectum (Tect)) for two brains using BARseq (Barcoded Anatomy Resolved by Sequencing, Chen et al., 2019), with $605$ and $704$ neurons observed for each brain. To summarize the MCMC samples of partitions from HBMAP, the $L=3$ particles of WASABI contain between 18 and 22 clusters (further details in Section E.1 of the Appendix). The clustering and projection patterns (referred to as projection motifs in the literature) identified by each particle are illustrated in Figure 7; dashed lines identify the clusters of neurons, and transparency reflects empirical projection strength (normalized barcode counts). To understand global differences between particles, neurons are colored by their group contribution to the VI distance between the first particle and each subsequent particle. We find that WASABI identifies a number of distinct projection motifs that are stable across the particles, particularly those characterizing projection to a single target area, such as the groups of neurons projecting only to Thal, Amyg, or Vislp. On the other hand, other motifs show more uncertainty, specifically, those characterizing more broad projections to two or more areas. For example, consider the group with high projection strength to AudC and low strength to Cstr and Amyg in particle 1. In particle 2, it is merged with another group of neurons with high strength to AudC and low strength to other areas in the ipsilateral cortex; instead, in particle 3, it is further split based on neurons with relatively lower strength to AudC. Another example is provided by neurons projecting to both Thal and Tect; these clusters are regrouped across the particles depending on their relative strength to these regions and if they exhibit low strength to Cstr. Overall, WASABI finds a number of distinct projection motifs from the auditory cortex, but some may have overlap or be further separated or merged, particularly those projecting more broadly across more than one target area.
Figure 7: Summarizing uncertainty in projection patterns from the auditory cortex. The empirical projection strength (computed as the neuron’s barcode counts divided by its total count) for each neuron (row) across the 11 target areas (columns) is illustrated by transparency. Each panel depicts the clustering of a particle, with neurons belonging to different clusters separated by dashed lines. Neurons are colored by their VIC by group comparing each particle to the first particle, to understand global differences in clustering across particles.
@@ 6 Discussion
Interpreting uncertainty characterized by the posterior distribution is challenging in complex, high-dimensional Bayesian models. As the posterior is intractable in such models, algorithms for approximate inference, such as MCMC, are employed, which return tens of thousands of approximate posterior draws. In this work, we focus on clustering, where the posterior is defined on the high-dimensional, partially-ordered space of partitions. In this non-standard setting, we propose a WASserstein Approximation for Bayesian cluserIng (WASABI) to summarize the posterior draws with a small number of weighted partition estimates. Through extensive simulations and analyses, we showcase its utility in practical cases when some clusters are not well separated or the model may be misspecified. Indeed, we demonstrate how focusing on a single clustering estimate can be misleading in such cases and how WASABI captures the different modes of clustering.
The Wasserstein approximation requires selecting a metric on the partition space. We focus on the variation of information (Meilă, 2007), also exploiting its properties to propose novel tools to assess uncertainty in the cluster assignment of individual points. However, WASABI can be extended to other choices, such as Binder’s loss, ARI, or generalizations (Dahl et al., 2022), or to more general settings, e.g. Buch et al. (2024) develop a Bayesian approach to clustering in the presence of noise.
An efficient algorithm, along with various tools to visualize and quantify uncertainty, is provided in the R package at https://github.com/cecilia-balocchi/WASABI. In practice, only a small number of iterations are typically required, with each one scaling linearly with the number of data points and MCMC samples. To reduce the cost of the latter, WASABI allows subsampling of the MCMC draws. For the former, future work will explore combining WASABI with distributed methods (Ni et al., 2020).
WASABI is useful for improved interpretation of the posterior, but could also be relevant for posterior compression in order to reduce both storage (saving only the WASABI particles and weights) and the computational cost of predictive inference. For example, density estimation can be expensive, depending on the size of the grid or number of new locations. For Example 1, density estimates are provided in Section D.3 of the Appendix, where the WASABI-based estimates significantly reduce the cost by a factor of $1,000$ compared to MCMC, at only a small increase in error.
Beyond clustering, the Wasserstein approximation could be beneficial for other Bayesian models. Future work will investigate settings, such as variable selection, trees, networks, and more.
@@ Appendix A Proof to Proposition 1
Let $\bm{w}^{*}=(w^{*}_{1},\ldots,w^{*}_{L})$ and $\bm{\rho}^{*}=\{\rho^{*}_{1},\ldots,\rho^{*}_{L}\}$ respectively denote the vector of weights and the set of atoms of $q^{*}(\cdot)=\sum_{\ell=1}^{L}w^{*}_{\ell}\delta_{\rho^{*}_{\ell}}(\cdot)$. To solve eq. (2) we need to find the optimal $\bm{w}^{*}$ and $\bm{\rho}^{*}$, but also for a given $q$ we need to find the optimal coupling $J^{*}(q)$ that minimizes $\sum_{\rho,\rho^{\prime}}d_{VI}(\rho,\rho^{\prime})J(\rho,\rho^{\prime})$ and that satisfies the marginal constraints.
@@ Appendix B Detailed algorithm and initialization strategies
[…]
  • Complete Linkage Initialization: Analogous to average linkage initialization, but with complete linkage.
  • k-means++ style Initialization (++++): Analogous to the k-means++ technique (Vassilvitskii and Arthur, 2006), this strategy promotes diversity among initial particles by iteratively choosing the next particle among the MCMC draws with probability proportional to its VI distance from the closest already chosen particle. This increases the likelihood that the $L$ initial particles will be well separated in partition space, which often leads to better local optima. The computational cost of this step is $O(TnL)$, corresponding to computing the VI distance between every MCMC draw and particle. Again thinning is applied by default to reduce the cost in $T$. Note that ++++ initialization does not require passing in or storing the posterior similarity matrix. In practice, this initialization requires multiple runs; however if the number of runs is sufficiently large, it often provides the best solution (smallest Wasserstein distance).
  • Lowest EVI Initialization (topvi): In this method, the expected VI is computed for all the partitions generated by the average linkage, complete linkage, and (if supplied) fixed initialization methods. The $L$ best partitions among these (with lower EVI) are then selected as initial centers. This approach combines average and complete linkage for a wider set of initial particles. The computational cost is similar to average or complete linkage initializaion (multiplied by a factor of two).
  • Fixed Initialization: This strategy allows the user to directly specify a set of $L$ fixed partitions to be used as initial centers. This is useful, for example, if the user wishes to include certain known clusterings (such as the minVI, MAP, Binder, or ARI estimators or expert-proposed solutions) as initial candidates.
A detailed description is provided in Algorithm 2. Note that the N-update step includes an optional outlier-search step, applicable if a region of attraction contains only one partition, indicating that the current center is not the closest to any other partition; this step tries to swap the center with a sampled “outlier” partition, i.e. a partition far (in VI sense) from all centers, and checks if the objective improves.
@@
   Update centers: $\rho^{*}_{\ell}\leftarrow\textsc{minVI}(\mathcal{N}_{\ell})$ $\triangleright$ Different search-algorithms can be used
@@
until change of $W$ is less than $\epsilon$ or $i>\texttt{max\_iter}$.
@@ Appendix C Additional material on describing and visualizing the particles
In Section 3.2, we introduce Example 2, where the data is generated from four Gaussian mixtures. We fit a diagonal location-scale Dirichlet Process Mixture model, using the R package BNPmix (Corradin et al., 2021), run for 10,000 iterations after a burn-in of 10,000 iterations. In the posterior similarity matrix shown in Figure 8, uncertainty about the number of clusters is apparent, but a subtle pattern of four equal-sized blocks can be observed, even though some of these seem to be often merged together.
@@
Figure 9 displays the posterior similarity matrices (PSMs) for the regions of attraction associated with each WASABI particle in Example 2. While Figure 3(b) (in Section 3.2) presents PSMs labeled by each particle’s own clusters (providing insight into the uncertainty structure around each specific clustering), here, the rows and columns are labeled according to the clusters defined by the meet of all particles. This common labeling facilitates direct comparison of cluster uncertainty across different regions of attraction, helping to identify points or groups whose allocation is most variable or most consistent between the particles represented by the WASABI summary.
@@ Appendix D Experiments: additional material
D.1 Slightly bimodal example
We consider the slightly bimodal example of Rajkowski (2019) with $n=600$ data points generated independently from a mixture of two Gaussians:
$Y_{i}\overset{\text{iid}}{\sim}0.5\text{N}(-1.1,1)+0.5\text{N}(1.1,1).$
Rajkowski (2019) focuses on the behavior of the MAP clustering when using a DP mixture of Gaussians and, in this example, shows that the partition with a single cluster has higher posterior mass than the intuitive partition of the data into two groups based on positive and negative observed values.
A Dirichlet process mixture of Gaussians is fit to data via the BNPmix package (Corradin et al., 2021). Specifically, the model assumes
$Y_{i}|P$ $\overset{\text{iid}}{\sim}\sum_{j=1}^{\infty}w_{j}\text{N}(\mu_{j},\sigma^{2}_{j}),$
$\mu_{j}|\sigma^{2}_{j}$ $\overset{\text{ind}}{\sim}\text{N}\left(\mu_{0},\frac{1}{k_{0}}\sigma^{2}_{j}\right),$
$\sigma^{2}_{j}$ $\overset{\text{iid}}{\sim}\text{IG}(a_{0},b_{0}),$
$w_{j}$ $=v_{j}\prod_{h<j}(1-v_{h}),\text{ with }v_{j}\overset{\text{iid}}{\sim}\text{Beta}(1,\alpha),$
with $\mu_{0}$ set to the sample mean, $k_{0}$ set to 0.1 to increase prior variability of the cluster centers, $a_{0}=3/2$ resulting in a marginal Student t prior on $\mu_{j}$ with $3$ degrees of freedom, and $b_{0}$ equal to the sample variance divided by 4. We consider two different choices of the DP parameter $\alpha$: 1) $\alpha=1$ which is a commonly-used setting, and 2) $\alpha=2/\log(n)$. The latter choice is motivated by Ascolani et al. (2023), who show that under regularity conditions, the marginal posterior of the number of clusters in the DPM can recover the truth, if $\alpha$ is tuned via an empirical or hierarchical approach.
DP parameter Median [95% CI] MAP VI VI.lb Binder ARI
$\alpha=1$ 7 [3,11] 1 1 2 59 4
$\alpha=2/\log(n)$ 4 [2,6] 1 1 2 2 2
Table 1: Slightly bimodal example. A comparison of the marginal posterior on the number of clusters (posterior median and 95% HPD credible interval) with the number of clusters in the MAP, VI, lowerbound to the VI (VI.lb), Binder, and ARI clustering estimates, for two choices of $\alpha=1$ and the empirical choice of $\alpha=2/\log(n)$.
(a) $\alpha=1$
(b) $\alpha=2/\log(n)$
Figure 12: Posterior similarity matrix for the slightly bimodal example for two choices of the DP parameter.
This example illustrates a paradox in BNP, where both overestimation and underestimation of the number of clusters occurs depending on how one choose to summarize the posterior. Specifically, it has been shown that the marginal posterior on the number of clusters can suffer from overestimation and inconsistency for the number of clusters if $\alpha$ is not properly tuned (Miller and Harrison, 2013, 2014; Ohn and Lin, 2023; Ascolani et al., 2023; Alamichel et al., 2024). In agreement with this, the 95% highest posterior density credible interval for the number of clusters shows overestimation when $\alpha=1$ in this slightly bimodal example, but contains the true number of clusters for the empirical choice of $\alpha$ (Table 1). Despite this, the optimal clustering estimate can still recover the truth under appropriate choice of loss (Rajkowski, 2019; Wade, 2023). However, in this example when the clusters are not well separated, the MAP and VI estimate contain only a single cluster, regardless of $\alpha$, while Binder and ARI overestimate the number of clusters for $\alpha=1$ and recover the true number for the empirical choice of $\alpha$ (Table 1). Interestingly, in previous literature, agreement was observed between the clustering solutions which minimize the expected VI and its more computationally efficient lower bound (Wade and Ghahramani, 2018); however, in this slightly bimodal example, they disagree, with the latter reporting two clusters (for both choices of $\alpha$). The posterior similarity matrix (Figure 12) shows greater uncertainty in clustering structure for $\alpha=1$.
(a) WASABI for $\alpha=1$
(b) WASABI for $\alpha=2/\log(n)$
Figure 13: Slightly bimodal example. WASABI’s particles are almost identical for both choices of $\alpha$ (the second particles differ in the allocation of only two points), but the weights are different, with higher mass on the one-cluster solution for the empirical choice of $\alpha$ and more equal weights for $\alpha=1$ in the WASABI posterior.
The WASABI posterior shows robustness to the choice of $\alpha$. For both choices, WASABI’s particles are almost identical, see Figure 13. However, the weights in the WASABI posterior are different, with higher mass on the one-cluster solution for the empirical choice of $\alpha$ (when the prior is more concentrated on a small number of clusters). Instead, the weights are more equal for $\alpha=1$.
D.2 Examples in Section 4